Accelerator Update for June 25, 2008 - June 27, 2008
Wednesday June 25
The day shift began with Operations monitoring the Tevatron (TeV) store 6246 (82.29E30) and the Recycler (RR) stash (214.1E10), with the Antiproton Source stacking (29.8E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
A controls expert reported at 9:44 AM that he found the front end (FE) for the MI flying wires had crashed.
Some time around 12:30 PM, Operations terminated store 6246.
Operations prepared D0 and CDF for access at 12:48 PM.
A water system expert reported at 1:06 PM that he was making up water for the TeV LCW system.
D0 came out of access at 1:27 PM.
CDF came out of access at 1:49 PM.
Operations put the TeV through a dry squeeze at 3:28 PM.
At 3:48 PM, Operations reported problems with the Recycler VSA. Operations contacted an expert.
Operations put the TeV into shot setup at 3:53 PM.
Operations reported at 4:01 PM that the Recycler had stashed antiprotons five times during the day shift.
At 4:18 PM, a Booster vacuum pump (B:PIRF23) tripped off. BRF13 & 14 also tripped off. Everything reset.
Operations reported at 4:27 PM that the Linac expert had fixed KRF3. BNB and Pbar resumed taking beam.
A controls expert reported at 4:38 PM that he had fixed the Recycler VSA.
Water experts reported at 6:01 PM that they were going to make up water for the MI RF and magnet LCW systems.
Operations established store 6248 at 6:13 PM with an initial luminosity of 280.9E30.
At 6:39 PM, LRF5 tripped off. It reset.
At 6:40 PM, Operations reported a problem with beam positions for store 6248. Operators adjusted the vertical tunes for both the protons and antiprotons.
At 8:33 PM, a P150 transfer line vacuum pump (IP7071) tripped off. It reset.
At 8:42 PM, Operations reported that they had been receiving many Booster watchdog trips, accompanied by BRF3 trips. BRF12 also seemed to spark about the same time as each trip, but operators were unsure if this was a cause or an effect. They turned off BRF12 as a test, but the trips continued. Operators brought on all the Booster RF stations and the number of trips dropped. Operators lowered the beam intensity for BNB to see if it would make a difference, but it didn’t. They brought BNB back to it normal intensity.
At 10:12 PM, Operations reported another storm of Booster watchdog trips. Operations contacted a Booster expert.
At 11:04 PM, LRF4 tripped off. It reset.
At 11:48 PM, there was another rash of watchdog trips. Operators reported that BRF12 was tripping “like crazy.” Operators turned off BRF12 and brought BRF19 on in its place.
At 11:52 PM, BRF18 tripped off and caused BRF19 to trip. Everything reset.
Operations reported at 11:59 PM that the Recycler had stashed four times during the evening shift.
Thursday June 26
The midnight shift began with Operations monitoring store 6248 (123.13E30) and the stash (159.2E10), with Pbar stacking (25.0E10), and with BNB taking beam.
At 12:12 AM, Booster suffered from another flurry of watchdog Anode Program High Limit trips over the next hour. BRF3 also tripped off several times.
Operations reported at 1:02 AM that they had lowered the Booster intensity to nine turns.
Around 2:44 AM, Operations brought BRF12 back on line and turned off BRF19.
At 2:46 AM, D0 reported that there was water dripping on their solenoid power supply. They turned off the power supply and rigged a tent to divert the water. They had the supply back on about an hour later.
Operations reported at 3:26 AM that they had completed some MI orbit smoothing.
Operations established beam to MTest at 4:43 AM.
At 5:01 AM, LRF4 & 5 tripped off. They both reset.
At 5:03 AM, KRF6 tripped off. It reset.
At 6:52 AM, a Booster vacuum pump (IPRF17) tripped off. BRF7 & 8 also tripped off. Everything reset.
Operations reported at 8:01 AM that the Recycler had stashed five times during the midnight shift.
At 8:16 AM, Operations reported more Booster watchdog trips.
At 9:28 AM, the Linac Klystron Buncher coil faulted. Experts made repairs and had it back online within 20 minutes.
Operations terminated store 6248 at 9:36 AM.
Operations put Pbar into standby at 9:45 AM.
A Booster expert turned off the Booster watchdog anode program at 10:01 AM. He had the program back on 15 minutes later.
Pbar resumed stacking at 10:57 AM.
TeV and MI experts began injection studies at 11:01 AM.
The injection studies ended and Operations put the TeV into shot setup at 12:41 PM.
The Water group reported at 2:36 PM that they were bringing the MI-50 heat exchanger back in to service.
Operations established store 6250 at 2:42 PM with an initial luminosity of 285.39E30.
At 3:19 PM, the MI High Level RF system tripped off during a stash and a transfer of antiprotons were lost. Operators investigated.
Operations put Pbar into standby at 3:26 PM.
At 3:31 PM, Operations reported that the Recycler 521 corrector magnet chassis was off. An expert investigated and brought it back online.
Operations reported at 3:49 PM that they were rebooting the MI Low Level RF system.
At 3:55 PM, a MI power supply (I:VT703) tripped off and caused the P1 transfer line beam permit to trip off. Everything reset.
Operations reported at 4:03 PM that the Recycler had stashed two times during the day shift.
At 4:12 PM, a Booster kicker (B:MP02) tripped off. It reset.
At 8:57 PM, Operations received a Pbar collection lens chilled water input pressure alarm. An expert investigated and concluded that a transducer had failed. Operators bypassed the alarm.
Operations disabled the MTest beam at 10:12 PM.
Operations reported at 11:57 PM that the Recycler had stashed five times during the evening shift.
Friday June 27
The midnight shift began with Operations monitoring store 6250 (94.57E30) and the stash (203.0E10), with Pbar stacking (36.5E10), and with BNB taking beam.
Operations conducted a Booster injection tune at 5:17 AM.
At 5:41 AM, KRF5 tripped off. It reset.
Operations established beam to MTest around 7 AM.
Operations terminated store 6250 at 7:49 AM.
Operations put the TeV into shot setup at 7:57 AM.
Operations reported at 8 AM that the Recycler had stashed six times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 06/30/2008 email Fermilab|
|Security, Privacy, Legal|